By the first principle of differentiation,find the derivative of f(X)=2/2x-1 at X=1


1
Expert's answer
2021-08-26T13:05:27-0400

Let us find the derivative of "f(x)=\frac{2}{2x-1}" at "x=1" using the first principle of differentiation:


"f'(1)=\lim\limits_{x\to 1}\frac{f(x)-f(1)}{x-1}
=\lim\limits_{x\to 1}\frac{\frac{2}{2x-1}-\frac{2}{2\cdot 1-1}}{x-1}
=\lim\limits_{x\to 1}\frac{\frac{2}{2x-1}-2}{x-1}
=\lim\limits_{x\to 1}\frac{\frac{2-4x+2}{2x-1}}{x-1}\\
=\lim\limits_{x\to 1}\frac{4-4x}{(2x-1)(x-1)}
=\lim\limits_{x\to 1}\frac{-4(x-1)}{(2x-1)(x-1)}
=\lim\limits_{x\to 1}\frac{-4}{2x-1}
=\frac{-4}{2\cdot 1-1}=-4."